# AReaL: A Large-Scale Asynchronous Reinforcement Learning System for Language Reasoning

### v1

@@ 7.2 End-to-End Comparison
We establish two state-of-the-art baselines using synchronous RL systems: DeepScaleR [26] for mathematical reasoning with a 1.5B model, and DeepCoder [25] for code generation with a 14B model, both trained using verl [46]. For larger 7B and 32B models where comparable baselines are unavailable, we performed controlled experiments by training from scratch using a synchronous variant of AReaL. After training, we evaluate on the challenging AIME24 benchmark for math models and the LiveCodeBench [15] benchmark for coding models. They are considered to the golden benchmarks and provide the most truthful reflection of reasoning capability improvement after RL, which was also used by models released by large companies [7, 32, 33, 54, 6]. Evaluation results on more benchmarks are presented in Appendix B.
Our main results are shown in Table 1. Since the code of obtaining previous SOTA models can be out-of-date, we measure the throughput and estimate the training hours using the latest verl code for a fair comparison. AReaL consistently matches or exceeds baseline performance while achieving significant speedups. In particular, our system demonstrates up to 2.57$\times$ improvement in training throughput compared to synchronous approaches without significant performance degradation.
Table 1: End-to-End Performance Comparison. We evaluate on the AIME24 benchmark for math and LiveCodeBench (8/1/24-2/1/25) for coding. We limit the maximum generation length to 32K tokens and sample 32 responses per question, reporting the average pass@1 accuracy. * represents the best known reproducible results obtained via RL , as cited from DeepScaler [26] and DeepCoder [25] respectively. AReaL achieve the comparable performance with 2x less training hours.
@@
w/ VeRL 43.1* 16 1000 33.6
w/ Sync.AReaL 42.0 16 1000 41.0
w/ AReaL (ours) 42.2 16 1000 14.8
@@
w/ VeRL - 24 1000 52.1
w/ Sync.AReaL 63.0 24 1000 57.7
w/ AReaL (ours) 63.1 24 1000 25.4
@@
w/ VeRL 57.9* 32 320 44.4
w/ Sync.AReaL 56.7 32 320 48.8
w/ AReaL (ours) 58.1 32 320 21.9
@@
w/ VeRL - 48 240 46.4
w/ Sync.AReaL 61.2 48 240 51.1
w/ AReaL (ours) 61.0 48 240 31.1
@@ 7.3 Scalability
We compare the scalability of AReaL with verl [46], the state-of-the-art synchronous RL system, across different model sizes and context lengths. AReaL’s throughput surpasses the baseline in most settings, and could achieve at most 2.57×2.57$\times$2.57 × speedup. We select the minimum number of GPUs when verl does not encounter the OOM issue for 7B models and 32k context length, then we proportionally adjust the number of GPUs according to the model size. We measure the effective throughput for training, defined as the rate of consuming generated tokens during PPO updates, after proper warmup steps. Figure 4 presents the results for context lengths of 16k and 32k. Here, context length refers to the sum of prompt length and generated length, with the maximum prompt length capped at 1k.
@@
Across all settings, AReaL demonstrates an approximate linear scaling trend with increased device count, while the synchronous system typically fails to scale effectively. We note that for smaller context lengths, the advantage of AReaL can be smaller because the generation throughput cannot match the pace of training throughput. Although many sequences are generated, they are not effectively consumed by the training process. Additionally, AReaL is more robust with longer generation lengths due to asynchronous and interruptible generation. The generation of long responses can be fully hidden in the critical path, so extending generation length does not drastically affect the effective training throughput of AReaL.
@@ 7.4 Algorithm Ablations
Figure 5: Ablation studies of the decoupled PPO objective and staleness control. Both algorithmic choices are essential. With a moderate staleness value and the decoupled objective, training progress can be accelerated by 2x while maintaining final evaluation performance.
Table 2: Evaluation scores when varying data staleness, comparing performance with and without the decoupled objective. Numbers within ±1 of the oracle score are underlined.
@@
We conduct ablation studies to validate our algorithmic innovations in Section 5 by training a 1.5B LRM on math tasks. Specifically, we vary the maximum allowed staleness $\eta$ and compare configurations with and without the decoupled PPO objective. Figures 5(b) and 5(a) show the learning curves after 1000 PPO updates. Table 2 presents the corresponding final evaluation performances across multiple mathematical reasoning benchmarks.

### v2

@@ 7.2 End-to-End Comparison
We establish two state-of-the-art baselines using synchronous RL systems: DeepScaleR [25] for mathematical reasoning with a 1.5B model, and DeepCoder [24] for code generation with a 14B model, both trained using verl [45]. For larger 7B and 32B models where comparable baselines are unavailable, we performed controlled experiments by training from scratch using a synchronous variant of AReaL. After training, we evaluate on the challenging AIME24 benchmark for math models and the LiveCodeBench [14] benchmark for coding models. Evaluation results on additional benchmarks are presented in Appendix B.
Our main results are shown in Table 1. Since the code for obtaining previous SOTA models can be out-of-date, we measure the throughput and estimate the training hours using the latest verl code for a fair comparison. AReaL consistently matches or exceeds baseline performance while achieving significant speedups without performance degradation. In particular, the end-to-end training time can be reduced by 2.77×2.$7$ compared with synchronous systems.
Table 1: End-to-End Performance Comparison. We evaluate on the AIME24 benchmark for math and LiveCodeBench (8/1/24-2/1/25) for coding. We limit the maximum generation length to 32K tokens and sample 32 responses per question, reporting the average pass@1 accuracy. * represents the best known reproducible results obtained via RL, as cited from DeepScaler [25] and DeepCoder [24] respectively. AReaL achieves comparable performance with 2$\times$ fewer training hours.
@@
w/ VeRL 43.1* 16 250 33.6
w/ Sync.AReaL 42.0 16 250 41.0
w/ AReaL (ours) 42.2 16 250 14.8
@@
w/ VeRL - 24 250 52.1
w/ Sync.AReaL 63.0 24 250 57.7
w/ AReaL (ours) 63.1 24 250 25.4
@@
w/ VeRL 57.9* 32 80 44.4
w/ Sync.AReaL 56.7 32 80 48.8
w/ AReaL (ours) 58.1 32 80 21.9
@@
w/ VeRL - 48 60 46.4
w/ Sync.AReaL 61.2 48 60 51.1
w/ AReaL (ours) 61.0 48 60 31.1
@@ 7.3 Scalability
We compare the scalability of AReaL with verl [45], the state-of-the-art synchronous RL system, across different model sizes and context lengths. We select the minimum number of GPUs when verl does not encounter the OOM issue for 7B models and 32k context length, then we proportionally adjust the number of GPUs according to the model size. We measure the effective throughput for training, defined as the rate of consuming generated tokens during PPO updates, after proper warmup steps. Figure 4 presents the results for context lengths of 16k and 32k. Here, context length refers to the sum of prompt length and generated length, with the maximum prompt length capped at 1k.
@@
Across all settings, AReaL demonstrates an approximate linear scaling trend with increased device count, while the synchronous system typically fails to scale effectively. AReaL’s throughput surpasses the baseline in most settings, and could achieve at most 2.5×2.5$\times$2.5 × speedup. We note that for smaller context lengths, the advantage of AReaL can be smaller because the generation throughput cannot match the pace of training throughput. Although many sequences are generated, they are not effectively consumed by the training process. Additionally, AReaL is more robust with longer generation lengths due to asynchronous and interruptible generation. The generation of long responses can be fully hidden in the critical path, so extending generation length does not drastically affect the effective training throughput of AReaL.
@@ 7.4 Algorithm Ablations
Figure 5: Ablation studies of the decoupled PPO objective and staleness control with a 1.5B model on math reasoning tasks. Both algorithmic choices are essential. With a moderate staleness value and the decoupled objective, training progress can be accelerated by over 2×2$\times$2 × while maintaining final evaluation performance.
Table 2: Evaluation scores when varying data staleness, comparing performance with and without the decoupled objective. Numbers within ±$\pm 1$ of the oracle score are underlined.
@@
We conduct ablation studies to validate our algorithmic innovations in Section 5 by training a 1.5B LRM on math tasks. We follow the basic experiment setting of DeepScaleR and then gradually increase the $\eta$ value for ablation purposes. Specifically, we vary the maximum allowed staleness $\eta$ and compare configurations with and without the decoupled PPO objective. Figures 5(b) and 5(a) show the learning curves after 250 training steps. Table 2 presents the corresponding final evaluation performances across multiple mathematical reasoning benchmarks. We follow the common practice of PPO and perform multiple mini-batch updates within each training step. We emphasize that $\eta$ constrains the training batch staleness regarding training steps.